OpenStudy (anonymous):

Find the sum and please show work. 11/12-5/16+11/18 (These are all fractions by the way...)

OpenStudy (sleepyjess):

Do you know what you need to do first?

OpenStudy (anonymous):

Find common denominators?

OpenStudy (anonymous):

Get them to all have the same denominator, 12*16*18=3456, then add or subtract and finally try to simplify the fraction you end up with.

OpenStudy (sleepyjess):

Yes

OpenStudy (sleepyjess):

Whatever you do to the denominator, you also have to do to the numerator

OpenStudy (anonymous):

Okay, so what would be the common denominators for all of them?

OpenStudy (sleepyjess):

douglas gave you a common denominator of 3456 by multiplying the numbers together

OpenStudy (sleepyjess):

Then you would need to take the 3456 and divide by 12, 16 and 18 so that you know what to multiply the numerator by

OpenStudy (anonymous):

I'm not sure that I understand... :(

OpenStudy (sleepyjess):

If you take 3456/12 = 288

OpenStudy (anonymous):

My final answer on wolframalpha.com was 1 31/144

OpenStudy (sleepyjess):

So you would take 11 * 288 = 3168

OpenStudy (anonymous):

4/12 becomes 20/60 by multiplying top and bottom by 5, for example

OpenStudy (sleepyjess):

Your first equation is 3168/3456

OpenStudy (sleepyjess):

First fraction

OpenStudy (sleepyjess):

Correct on your 4/12

OpenStudy (sleepyjess):

2nd fraction of 5/16

OpenStudy (anonymous):

Could you please write this out for me? I do not understand, and I am more of a visual learner.

OpenStudy (sleepyjess):

3456/16 = 216 216 * 5 = 1080 so your 2nd fraction would be 1080/3456

OpenStudy (sleepyjess):

3rd fraction of 11/18 3456/18 = 192 192 * 11 = 2112/3456

OpenStudy (sleepyjess):

So now lets put those all together

OpenStudy (sleepyjess):

3168/3456 - 1080/3456+2112/3456

OpenStudy (sleepyjess):

Now add (or subtract) the numerators

OpenStudy (anonymous):

I'm lost, I'm sorry, I think I'll just skip this problem.

OpenStudy (anonymous):

Thank you very much for your help anyways-I'll give you a medal for your awesomeness. :)

OpenStudy (sleepyjess):

Do you understand how to get the common denominator?

OpenStudy (sleepyjess):

I want you to understand it though... it will help you going forward

OpenStudy (anonymous):

Yes, I understand the common denominator, I just don't understand combining them all at once...

OpenStudy (sleepyjess):

When you have the common denominator, you have to figure out how many times the original denominator goes into that number, so that you can multiply the numerator by that number... each one by itself \[\frac{ 11 }{ 12 }-\frac{ 5 }{ 16 }+\frac{ 11 }{ 18 }\] Common Denominator = 12*16*18 = 3456 So on \[\frac{ 11 }{ 12 }\] \[\frac{ 3456 }{ 12 }=288\]Then you take the 11 (the numerator) and multiply it by 288 \[\frac{ 11*288 }{ 3456 }=\frac{ 3168 }{ 3456 }\]

OpenStudy (sleepyjess):

Do you understand that part?

OpenStudy (anonymous):

I don't understand the second step.

OpenStudy (sleepyjess):

Okay... we determined that 3456 was a common denominator of all 3 numbers

OpenStudy (sleepyjess):

So we know what our number is going to be for each of the fractions on the denominator, but remember we have to do the same thing to the numerator as we do for the denominator. The first fraction... in order to get 3456 on the bottom it would be 12*288 = 3456 (12 being the original number)... in order to get the 288, all I did was divide 3456 by 12

OpenStudy (anonymous):

Thanks, but I have to go. I am going to look over this later and see if I can solve it. Thank you very much, you are amazing! :D

OpenStudy (sleepyjess):

Okay... I'm attaching a file for you that might make it easier.

OpenStudy (anonymous):

Nicely done.